OpenStudy (anonymous):

Which of the following forms a proportion with ten over fourteen? twenty-one over fifteen two over seven twenty-five over thirty-five seven over five

OpenStudy (anonymous):

the answer is C

OpenStudy (anonymous):

Do u wanna know why?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first step simplest form... 10/2=5.....14/2=7

OpenStudy (anonymous):

so now you have to find wich answer comes to 5over7

OpenStudy (anonymous):

25/5=5.....35/5=7.....5over7

OpenStudy (anonymous):

U follow?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so (C) is the answer

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

anytime :)